------------------------------------------------------------ Date: Wed Feb 25 07:54:18 2004, Shift: 00:00-08:00 Supervisor: EBJ, with: Log book#15 pages 67 - 67 Shift Summary: ------------------------------------------------------------ 1. We had good beam until 6am, when they began beam study. 2. I took a lot of data for the MRS at full field and the FS at 8deg and at 1/2 field. 3. The detector had no problems. ------------------------------------------------------------ _______________________________________________ Brahms-l mailing list Brahms-l@lists.bnl.gov http://lists.bnl.gov/mailman/listinfo/brahms-lReceived on Wed Feb 25 07:54:34 2004
This archive was generated by hypermail 2.1.8 : Wed Feb 25 2004 - 07:54:54 EST